# Relativistic structure of a supermassive black hole embedded in the dark matter halo of NGC 4649 (M60)

### v1

@@ Abstract
We construct a static, spherically symmetric black hole (BH) solution embedded within a dark matter (DM) halo, formulated as a non-vacuum extension of the Schwarzschild spacetime. The DM distribution is modeled via an empirical density profile calibrated to observations of the elliptical galaxy NGC 4649 (M60), incorporating Hubble Space Telescope (HST) imaging, stellar velocity dispersion data, and globular cluster dynamics. The resultant spacetime metric depends on three independent parameters: the black hole mass $M$, the asymptotic circular velocity $V_{c}$, and the halo scale radius $a$, and smoothly reduces to the Schwarzschild limit as $V_{c}\to 0$ and $a\to 0$. We analyze the influence of the halo on key geometric and physical quantities, including the event horizon radius, photon sphere, shadow size, and curvature invariants. The Kretschmann scalar exhibits an enhanced sensitivity to halo-induced modifications, particularly in the near-horizon regime. Thermodynamic properties of the solution are also examined. In the extremal limit, characterized by a vanishing surface gravity, the model supports a finite tangential pressure, implying a non-trivial extension of standard black hole thermodynamics. These results highlight the relevance of incorporating astrophysical environments into BH modeling and offer new avenues for testing strong-field gravity through precision observational data.
@@ I Introduction
This recognition has motivated a growing body of research into how BHs interact with surrounding DM halos. Unlike idealized models of isolated BHs in asymptotically flat spacetimes, astrophysical BHs reside in complex galactic environments shaped by DM [27, 26, 25]. Recent theoretical and numerical studies show that such environments can significantly influence BH behavior. In particular, DM may affect accretion dynamics, perturb horizon geometry, and modify key thermodynamic quantities [29, 28, 30, 31]. These interactions may also impact BH growth and feedback over cosmological timescales. Collectively, these findings highlight the need to account for DM in any realistic model of BH physics, especially when interpreting observations or testing gravity in strong-field regimes.
Moreover, DM density profiles, including the widely used Navarro-Frenk-White (NFW) model [32], along with its various alternatives [34, 35, 33], have played a vital role in modeling the gravitational potential on galactic scales. These profiles provide crucial input for understanding the structure and dynamics of galaxies. When embedded into relativistic spacetime geometries, such profiles enable the construction of BH metrics that exhibit significant departures from the standard Schwarzschild or Kerr solutions [36, 37, 39]. These generalized configurations offer a powerful framework for studying BHs not only as high-energy astrophysical engines, but also as precision probes of DM distributions and their underlying physical properties [41, 40, 42].
@@ II.1 Spacetime metric
We consider the spacetime line element given by
@@
where the angular surface element in spherical coordinates is given by $d\Omega^{2}=d\theta^{2}+\sin^{2}{\theta}\,d\phi^{2}$. In the general case, $f(r)$ and $g(r)$ are distinct functions.
@@ II.3 Specific Solution
$G_{\mu\nu}=8\pi T_{\mu\nu}^{\rm(DM)}\ ,$ (5)
where $\left(T^{\rm(DM)}\right)^{\mu}_{\ \nu}={\rm diag}(-\rho,0,P,P)$. According to the definition of the density profile $\rho=\rho(r)$, the mass function $M(r)$, the metric coefficient $F(r)$ and the pressure $P(r)$ can be determined by solving the following coupled field equations:
$M^{\prime}(r)=4\pi r^{2}\rho(r)\ ,$ (6)
$\frac{F^{\prime}(r)}{F(r)}=\frac{2M(r)}{r[r-2M(r)]}$ (7)
and
$P(r)=\frac{M(r)\rho(r)}{2[r-2M(r)]}.$ (8)
Substituting the density (1) in Eq. (6), leads to the mass function
@@ III Geometric and Thermodynamic Properties
Table 1: Parameters $V_{c}$, $a$ and $M$ in different units with minimum and maximum values.
@@ III.1 Event Horizon
According to Eq. (18), this model determines a low critical tangential velocity for the orbital radius of the halo, which agrees with the data of [36]. This shows that high velocities can lead to instability. However, Eq. (19) shows that with increasing $V_{c}$ the range of permissible values for the radius $a$ becomes smaller, but to a greater extent.
Figure 1 shows that the constraints on $a$ are more restrictive for more massive BHs, i.e. the allowable range for the amount of $a$ becomes narrower at higher magnitudes. This behavior is not only consistent with the data in [36], but is also expected for larger BHs forcing more distant radii. For Data I and II we have: $a_{I}>5.96\times 10^{12}$, and $\xi_{I}-1=1.26\times 10^{-20}$; and $a_{II}>7.67\times 10^{12}$, and $\xi_{II}-1=1.67\times 10^{-20}$.
Figure 2 shows the behavior of $F(r)$, related to the temporal component of Eq. (15), and shows slight changes in the values of $r_{h}$ for each data set. For Data I, Data II and without halo, respectively: $r_{hI}=1.034\times 10^{13}\textnormal{m}$, $r_{hII}=1.33\times 10^{13}\textnormal{m}$ and $r_{h}^{\textnormal{(sch)}}=1.27\times 10^{13}\textnormal{m}$. The presence of the DM halo has a certain influence on $r_{h}$, which indicates that for higher values of the parameters $(m_{BH},V_{c},a)$ there is an increase in the radius compared to the case without halo. For lower values of the same parameters, the radius decreases.
@@ III.2 Kretschmann Scalar
$+\frac{8\left[m_{BH}\left(a^{2}+r^{2}\right)^{2}-2a^{2}r^{3}V^{2}\right]^{2}}{r^{6}\left(a^{2}+r^{2}\right)^{4}}+\frac{1}{8r^{6}}\left(\frac{2[\lambda+(\xi-1)r]\left[m_{BH}(a^{2}+r^{2})^{2}-2a^{2}r^{3}V^{2}\right]}{(r-\lambda)\left(a^{2}+r^{2}\right)^{2}}+\frac{r\tau_{3}(r)}{g(r)(r-\lambda)^{2}}\right)^{2},$
@@
$\tau_{1}(r)=2m_{BH}\left(4\lambda^{2}+\left(\xi^{2}-5\xi+4\right)r^{2}+\lambda(3\xi-8)r\right),$ (21)
$\tau_{2}(r)/r$ $=$ $(\xi-1)r^{2}\left(-\xi+(\xi-1)V^{2}+3\right)$ (22)
$-2\lambda r\left(\xi+V^{2}-3\right)+\lambda^{2}\left(V^{2}-3\right),$
$\tau_{3}(r)$ $=$ $(\lambda+(\xi-1)r)^{2}-2r(\lambda+(\xi-1)r)$ (23)
$-2(r-\lambda)(\lambda+(\xi-1)r)+2(\xi-1)r(r-\lambda),$
@@
$\tau_{4}(r)=r\left(3\lambda^{2}+\left(\xi^{2}-4\xi+3\right)r^{2}+2\lambda(\xi-3)r\right).$ (24)
@@
$\lim_{r\to\infty}K(r)\to 0\,,\qquad{\rm and}\qquad\lim_{r\to 0}K(r)\to\infty$ (25)
@@
Figure 3: Graphical representation of the scalar $K(r)$ from Eq. (20). For the value sets: ($V_{cI},a_{I},M_{I}$), given by Data I, ($V_{cII},a_{II},M_{II}$), given by Data II, and for the Schwarzschild-type solution without a halo (0,0,$M_{0}$).
Analogous to the case without a halo, Fig. 3 shows that for higher values of the parameters $(m_{BH},V_{c},a)$, which correspond to Data II, the values of $(K(r),r)$ are slightly larger. For Data I, on the other hand, the lower values of the same parameters indicate a slight decrease in curvature, indicated by a shift to the left. This suggests a relationship between mass and curvature, where BHs with larger mass, which is increased by the halo, exhibit curvature that extends over a larger surrounding area. The reverse behavior is observed for smaller masses.

### v2

@@ Abstract
We construct a static, spherically symmetric black hole (BH) solution embedded within a dark matter (DM) halo, formulated as a non-vacuum extension of the Schwarzschild spacetime. The DM distribution is modeled via an empirical density profile calibrated to observations of the elliptical galaxy NGC 4649 (M60), incorporating Hubble Space Telescope (HST) imaging, stellar velocity dispersion data, and globular cluster dynamics. The resultant spacetime metric depends on three independent parameters: the black hole mass $M$, the asymptotic circular velocity $V_{c}$, and the halo scale radius $a$, and smoothly reduces to the Schwarzschild limit as $V_{c}\to 0$ and $a\to 0$. We analyse the influence of the halo on key geometric and physical quantities, including the event horizon radius, photon sphere, shadow size, and curvature invariants. The Kretschmann scalar exhibits an enhanced sensitivity to halo-induced modifications, particularly in the near-horizon regime. Thermodynamic properties of the solution are also examined. In the extremal limit, characterized by a vanishing surface gravity, the model supports a finite tangential pressure, implying a non-trivial extension of standard black hole thermodynamics. These results highlight the relevance of incorporating astrophysical environments into BH modelling and offer new avenues for testing strong-field gravity through precision observational data.
@@ I Introduction
This recognition has motivated a growing body of research into how BHs interact with surrounding DM halos. Unlike idealized models of isolated BHs in asymptotically flat spacetimes, astrophysical BHs reside in complex galactic environments shaped by DM [27, 26, 25]. Recent theoretical and numerical studies show that such environments can significantly influence BH behaviour. In particular, DM may affect accretion dynamics, perturb the horizon geometry, and modify key thermodynamic quantities [29, 28, 30, 31]. These interactions may also impact BH growth and feedback over cosmological timescales. Collectively, these findings highlight the need to account for DM in any realistic model of BH physics, especially when interpreting observations or testing gravity in strong-field regimes.
Moreover, DM density profiles, including the widely used Navarro-Frenk-White (NFW) model [32], along with its various alternatives [34, 35, 33], have played a vital role in modelling the gravitational potential on galactic scales. These profiles provide crucial input for understanding the structure and dynamics of galaxies. When embedded into relativistic spacetime geometries, such profiles enable the construction of BH metrics that exhibit significant departures from the standard Schwarzschild or Kerr solutions [36, 37, 39]. These generalized configurations offer a powerful framework for studying BHs not only as high-energy astrophysical engines, but also as precision probes of DM distributions and their underlying physical properties [41, 40, 42].
@@
In our analysis, we chose not to adopt the traditional NFW profile. Although this model provides a realistic fit to the kinematics of globular clusters (GCs), it lacks the necessary central concentration to account for mass dominance in the vicinity of the BH [36, 48]. While the NFW profile does exhibit a rising density toward the center, this increase is insufficient to accurately represent the central mass distribution. Alternatively, the density profile $\rho_{\rm DM}$, given by Eq. (1), not only provides a statistically robust fit to the data from the M60 galaxy [36], supporting analyses involving the parameters $a$ and $V_{c}$, but also exhibits a stable, non-divergent behavior as $r\to 0$. This indicates the presence of a central core rather than a cusp, a feature that aligns more closely with observational data. The regular behavior at the center is a desirable property in density models, as it enhances consistency with empirical observations [35].
@@ II.1 Spacetime metric
In the context of static and spherically symmetric solutions, it is generally appropriate to consider that $g_{tt}\neq-1/{g_{rr}}$, as the equality arises only in specific cases [49]. Under this more general assumption, the line element can be expressed in the following form:
@@
where the angular surface element in spherical coordinates is given by $d\Omega^{2}=d\theta^{2}+\sin^{2}{\theta}\,d\phi^{2}$. In the general case, $F(r)$ and $G(r)$ are functions of the radial coordinate. The Schwarzschild metric represents a specific case in which the metric functions satisfy $F(r)=G(r)=1-2m_{\rm BH}/r$, where $m_{\rm BH}$ denotes the constant mass of the BH [1]. By relaxing the condition $F(r)=G(r)$, one obtains generalized extensions of the Schwarzschild solution. In the following sections, we will construct a solution where these functions differ.
In metric solutions describing static and asymptotically flat spacetimes, event horizons correspond to non-singular physical surfaces [50]. These are characterized by the vanishing of the temporal component of the metric [^1] In an asymptotically flat spacetime, the surface defined by the vanishing of the normalized square of the temporal Killing vector $\xi^{\mu}$, that is, $\xi^{\mu}\xi_{\mu}=g_{tt}=0$, constitutes a null hypersurface. This surface functions as an event horizon, beyond which no future-directed causal curve, tangent to null geodesics confined within that region, can escape [50]., i.e.,
$g_{tt}\to 0\ .$ (3)
This occurs because the Killing vector becomes null at the horizon, aligning with the direction of null geodesics and remaining tangent to them, thereby confined to the event horizon surface itself [51].
@@ II.3 Specific Solution
The spherically symmetric spacetime described in Eq. (2) can also be expressed in the form given by Eq. (5), where the functions $F(r)$ and $M(r)$ are introduced explicitly. By setting $F(r)=1-2m_{\rm BH}/r$ and $M(r)=m_{\rm BH}$, Eq. (5) reduces to the Schwarzschild solution. In what follows, we derive a Schwarzschild-type solution based on the density profile given in Eq. (1), which incorporates the presence of a dark matter (DM) halo. For specific parameter choices, this solution reduces to the standard Schwarzschild form.
@@
$G^{\mu}_{\ \nu}=8\pi\left(T^{\rm(DM)}\right)^{\mu}_{\ \nu}\,,$ (6)
where $\left(T^{\rm(DM)}\right)^{\mu}_{\ \nu}={\rm diag}(-\rho_{\rm DM},0,P,P)$, and $P$ is the tangential pressure. The relationships involving the density profile $\rho_{\rm DM}=\rho_{\rm DM}(r)$, the mass function $M(r)$, the metric coefficient $F(r)$ and the tangential pressure function $P=P(r)$ are found from the components of the energy-momentum tensor. Thus, setting $\left(T^{\rm(DM)}\right)^{t}_{\ t}=-\rho_{\rm DM}(r)$, we have
$M^{\prime}(r)=4\pi r^{2}\rho_{\rm DM}(r)\,.$ (7)
For $\left(T^{\rm(DM)}\right)^{r}_{\ r}=0$, where we consider the radial pressure to be zero due to the absence of radial motion of the matter constituting the halo, we have
$\frac{F^{\prime}(r)}{F(r)}=\frac{2M(r)}{r[r-2M(r)]}\,.$ (8)
Finally, for $\left(T^{\rm(DM)}\right)^{\theta}_{\ \theta}=\left(T^{\rm(DM)}\right)^{\phi}_{\ \phi}=P(r)$, where using Eqs. (7) and (8), we obtain
$P(r)=\frac{M(r)\rho_{\rm DM}(r)}{2[r-2M(r)]}.$ (9)
By solving the coupled field equations (7)–(9), we obtain the functions $M(r)$, $F(r)$ and $P(r)$.
Substituting the energy density (1) in Eq. (7), leads to the mass function
@@ III Geometric and Thermodynamic Properties
Table 1: Parameters $V_{c}$, $a$ and $M$ in different units with minimum and maximum values. The IS values were taken from [36].
@@ III.1 Event Horizon
As shown in Eq. (19), this model predicts a low critical tangential velocity for the orbital radius of the halo, consistent with the observational data reported in [36]. This indicates that higher velocities may lead to dynamical instability. Furthermore, Eq. (20) demonstrates that as $V_{c}$ increases, the range of admissible values for the radius $a$ decreases, though this reduction becomes progressively more significant.
Figure 1 illustrates that the constraints on the parameter $a$ become more restrictive for more massive black holes (BHs), meaning that the permissible range for $a$ narrows as the BH mass increases. This behavior is not only consistent with the observational data presented in [36], but is also physically expected, as more massive BHs tend to influence matter at larger characteristic radii. For Data I and II we have: $a_{I}>5.96\times 10^{12}$, and $\xi_{I}-1=1.26\times 10^{-20}$; and $a_{II}>7.67\times 10^{12}$, and $\xi_{II}-1=1.67\times 10^{-20}$.
Figure 2 shows the behavior of $F(r)$, related to the temporal component of Eq. (16), and shows slight changes in the values of $r_{h}$ for each data set. For Data I, Data II and without halo, respectively: $r_{hI}=1.034\times 10^{13}\textnormal{m}$, $r_{hII}=1.33\times 10^{13}\textnormal{m}$ and $r_{h}^{\textnormal{(sch)}}=1.27\times 10^{13}\textnormal{m}$. The radii $r_{hI}$ and $r_{hII}$ were computed using Eq. (18), based on the parameters from Data I and Data II, as specified above Table 1. It is important to note that, in the case without a halo, corresponding to the standard Schwarzschild solution, we employed the mass value $M_{0}$, defined just below Table 1. The presence of the dark matter (DM) halo influences the value of $r_{h}$, indicating that higher values of the parameters $(m_{\rm BH},V_{c},a)$ lead to an increase in the radius compared to the halo-free case. Conversely, for lower values of these parameters, the radius decreases. The variation in the horizon radius can be related to changes in mass. Specifically, the inequalities $M_{I}<M_{0}<M_{II}$ and $r_{hI}<r_{h}^{\text{(sch)}}<r_{hII}$ indicate that, for a spherical object with constant matter density, an increase in mass corresponds to an increase in its total surface area. This observation leads us to consider the influence of the halo on the black hole entropy $S$, which will be addressed in Subsection III.4.
@@ III.2 Kretschmann Scalar
$+\frac{8\left[m_{BH}\left(a^{2}+r^{2}\right)^{2}-2a^{2}r^{3}V_{c}^{2}\right]^{2}}{r^{6}\left(a^{2}+r^{2}\right)^{4}}+\frac{1}{8r^{6}}\left(\frac{2[\lambda+(\xi-1)r]\left[m_{BH}(a^{2}+r^{2})^{2}-2a^{2}r^{3}V_{c}^{2}\right]}{(r-\lambda)\left(a^{2}+r^{2}\right)^{2}}+\frac{r\tau_{3}(r)}{g(r)(r-\lambda)^{2}}\right)^{2},$
@@
$\tau_{1}(r)=2m_{BH}\left[4\lambda^{2}+\left(\xi^{2}-5\xi+4\right)r^{2}+\lambda(3\xi-8)r\right],$ (22)
$\tau_{2}(r)/r$ $=$ $(\xi-1)r^{2}\left[-\xi+(\xi-1)V_{c}^{2}+3\right]$ (23)
$-2\lambda r\left(\xi+V_{c}^{2}-3\right)+\lambda^{2}\left(V_{c}^{2}-3\right),$
$\tau_{3}(r)$ $=$ $[\lambda+(\xi-1)r]^{2}-2r[\lambda+(\xi-1)r]$ (24)
$-2(r-\lambda)[\lambda+(\xi-1)r]+2(\xi-1)r(r-\lambda),$
@@
$\tau_{4}(r)=r\left[3\lambda^{2}+\left(\xi^{2}-4\xi+3\right)r^{2}+2\lambda(\xi-3)r\right].$ (25)
Expanding Eq. (21) as a Taylor series, and considering the behavior of $K(r)$ for large values of $r$, we obtain the following expressions:
$K(r)=\frac{1}{r^{4}}\Bigg{[}\left(\xi-1\right)^{2}\left(\xi^{2}-6\xi+17\right)\left(\frac{1}{4}-V_{c}^{2}\right)$ (26)
$+V_{c}^{4}\left(\xi^{4}-8\xi^{3}+30\xi^{2}-40\xi+33\right)\Bigg{]}+\mathcal{O}\left(\frac{1}{r^{5}}\right),$
and when $r$ is very small, we have the following approximation:
$K(r)=\frac{48m_{BH}^{2}}{r^{6}}+\frac{20m_{BH}}{r^{5}}\left(\frac{2m_{BH}\xi}{\lambda}-1\right)+\mathcal{O}\left(\frac{1}{r^{4}}\right)\,.$ (27)
For the Schwarzschild solution, we have the scalar $K_{0}(r)$,
$K_{0}(r)=\frac{48m^{2}_{BH}}{r^{6}}\ .$ (28)
For regions sufficiently far from the black hole center, Eq. (26) shows that $K(r)\sim\frac{1}{r^{4}}$, indicating that the curvature decreases with increasing radius, as expected for asymptotically flat spacetimes. This implies that the halo’s contribution becomes negligible in this regime, vanishing regardless of the halo parameter values as $r\to\infty$. This behavior is confirmed by the main plot in Fig. 3.
In contrast, for regions near the black hole center, Eq. (27) and the inset in Fig. 3 show that the dominant term behaves as $\sim 1/r^{6}$, characteristic of the Schwarzschild solution. This indicates that the central singularity is not eliminated by the presence of the halo. Nonetheless, the halo contributes through subleading terms such as $\sim 1/r^{5}$ and others of lower order, primarily influenced by the parameters $V_{c}$ and $a$.
@@
$\lim_{r\to\infty}K(r)\to 0\,,\qquad{\rm and}\qquad\lim_{r\to 0}K(r)\to\infty\,,$ (29)
@@
Figure 3: Graphical representation of the scalar $K(r)$ from Eq. (21). For the value sets: ($V_{cI},a_{I},M_{I}$), given by Data I, ($V_{cII},a_{II},M_{II}$), given by Data II, and for the Schwarzschild-type solution without a halo (0,0,$M_{0}$). These values are listed immediately before and after Table 1. The main graph shows the behaviour for larger scales of $r$, and the inner graph refers to smaller scales of $r$.
Analogous to the case without a halo, Fig. 3 shows that for higher values of the parameters $(m_{BH},V_{c},a)$, which correspond to Data II, the values of $(K(r),r)$ are slightly larger. For Data I, on the other hand, the lower values of the same parameters indicate a slight decrease in curvature, indicated by a shift to the left. This suggests a relationship between mass and curvature, where BHs with larger mass, which is increased by the halo, exhibit curvature that extends over a larger surrounding area. The reverse behaviour is observed for smaller masses.
The overall behavior of $K(r)$ is not only illustrated in Fig. 3, but also presented numerically in Table 2, which highlights the differential behavior at scales near the black hole center and the transition toward the characteristic radius $a$ of the halo. Around $r\sim a$, the curvature exhibits a slower decay in the presence of the halo compared to the halo-free case. This indicates that, while the datasets reflect distinct curvature behavior near the central region, they also reveal a pronounced increase in curvature extending into the vicinity of the halo’s critical radius. In essence, the halo prolongs the persistence of the spacetime curvature, although this effect vanishes asymptotically, as described by Eq. (29).
$r$[m] $K_{0}(r)/K_{I}(r)$ $K_{0}(r)/K_{II}(r)$
$10^{-7}r_{h}^{\textnormal{(sch)}}=1.27\times 10^{6}$ $1.50857$ $0.91181$
$r_{h}^{\textnormal{(sch)}}=1.27\times 10^{13}$ $1.50857$ $0.911804$
$4\times 10^{6}r_{h}^{\textnormal{(sch)}}=5.08\times 10^{19}$ $1.2208$ $0.832969$
$0.19a_{I}=5.8634\times 10^{19}$ $0.980184$ $0.747298$
$a_{I}=3.086\times 10^{20}$ $0.000955153$ $0.000599851$
$a_{II}=4.629\times 10^{20}$ $0.000279914$ $0.000132006$
$2a_{II}=9.258\times 10^{20}$ $0.0000484906$ $0.0000179377$
Table 2: Ratio between the scalar $K_{0}(r)$, given by Eq. (28) for the Schwarzschild solution without a halo, and the scalar $K(r)$ corresponding to the halo-influenced solution, defined in Eq. (21). The functions $K_{I}(r)$ and $K_{II}(r)$ are derived from Data I and Data II, respectively.
Examining Table 2, particularly the first three rows corresponding to $r\sim r_{h}^{\text{(sch)}}$, we observe an increase in curvature intensity for Data II, in contrast to the attenuation observed for Data I. This behavior highlights the influence of the parameters $V_{c}$ and $a$ at scales on the order of $r_{h}^{\text{(sch)}}$, given that $(V_{cII},a_{II})>(V_{cI},a_{I})$. Moreover, the inset of Fig. 3 shows that even near the black hole center, where the curvature diverges, the scalar $K(r)$ remains sensitive to the halo parameters. Accordingly, considering the dark matter density profile $\rho_{\rm DM}$ given in Eq. (1), such an influence is expected to manifest prominently within the halo’s central core region.
Figure 4 presents the ratio $K_{0}(r)/K(r)$, previously discussed in Table 2, illustrating how the parameters $V_{c}$ and $a$ contribute to an increase in curvature as one approaches the black hole center. This ratio stabilizes within the halo’s “core” region, where it reaches the constant value $K_{0}(r)/K(r)=M_{0}/m_{\rm BH}$. As shown in Table 2, we find $K_{0}(r)=1.50857\,K_{I}(r)$ for Data I and $K_{0}(r)=0.91181\,K_{II}(r)$ for Data II. These results indicate that larger values of the halo parameters—i.e., $(V_{cII},a_{II})>(V_{cI},a_{I})$—lead to a reduction in the curvature within the halo’s core.
Figure 4: Graphical representation of the ratio $K_{0}(r)/K(r)$, where $K(r)$ corresponds to the solution with a DM halo as given in Eq. (21), and $K_{0}(r)$ represents the Schwarzschild solution without a halo from Eq. (28). The plots are generated using the parameter sets $(V_{cI},a_{I},M_{I})$ and $(V_{cII},a_{II},M_{II})$, corresponding to Data I and Data II, respectively. These values are detailed immediately before and after Table 1.